Use the formula V = 15/22√gd to determine the depth of the water for tsunami wave starveling at the given speeds. (V is the wave's speed in miles per hour, g is the gravitational acceleration of 32 ft/sec^2, and d is the depth of the water in feet.)
QUESTION:
180 mi/hr
NOTE:
√gd means the square root of gd

Answer by stanbon(75887)

Use the formula V = 15/22√gd to determine the depth of the water for tsunami wave starveling at the given speeds. (V is the wave's speed in miles per hour, g is the gravitational acceleration of 32 ft/sec^2, and d is the depth of the water in feet.)
QUESTION:
180 mi/hr
---------
180 = (15/22)sqrt(32*d)
32*22 = sqrt(32d)
----
32d = 704^2
---
d = 15488 ft
--------------
Cheers,
Stan H.
-------------